# NTSEBench: Cognitive Reasoning Benchmark for Vision Language Models

### v2

@@ 4 Results and Analysis
NTSEBench presents a challenging task for SOTA LLMs and VLMs. Based on the findings in Tables 2 and 3, it is evident that the proposed dataset presents a challenging task for all state-of-the-art LLM and VLM models. None of the open-source models achieve accuracy exceeding 50% on text-only questions and 35% on multi-modal questions, with propriety models achieving 62% and 42% accuracy respectively. Although o1-preview has created its own category and got more than 80% accuracy on text only data, it can not do multi-modal reasoning. Many of the models tested didn’t even reach random selection/guess baseline of 24.52%(261 question had 5 options and 2467 question has 4 options), indicating that current LLMs and VLMs struggle with cognitive reasoning questions.
@@
Options Bias Ablation. From Gemini 1.5 Pro Results in Table  8 and  9 we can see that aggregate performance for both multi-modal and text question is affected by the position of the correct option(variation in performance ranging from -4 to +6 percent when compared to random for text and -5 to +5 percent for multi-modal categories). We can make similar observation on GPT-4o (tables  10 and  11), however, GPT-4o has smaller variation in performance than Gemini Pro. These variations suggest that the model’s responses to logical questions may be either memorized or guessed, indicating a potential limitation in LLMs and VLMs. Even o1-preview, as shown in Table 12, exhibits performance variations akin to those of GPT-4o, suggesting that advanced reasoning models are similarly affected by bias.
Error Analysis. We manually conducted an error analysis of 260 questions (10 from each question category) for Gemini 1.5 pro, and we identified distinct patterns in reasoning and error categorization. We have categorised errors based on the cognitive dimensions outlined in section 2. The Sankey diagram in figure 2 illustrates how errors across various question categories correspond to specific error types. A key observation is that many errors arise from Pattern Recognition failures, especially in categories like Alphabet Tests, Non-Verbal Analogy, and Series questions, where the model struggled with recurring patterns and sequence shifts, highlighting challenges in complex pattern-based reasoning. We also noted frequent errors in Spatial Reasoning and Logical Deduction tasks, particularly in spatial or diagrammatic questions like Cube and Dice, Embedded Figure, and Paper Folding & Cutting. These questions often require pattern recognition, shape manipulation, or deducing logical relations from limited visual data. The figure shows that errors in Quantitative Analysis were common in numerical questions like Time and Clock and Mathematical Operations, indicating the model excels in simpler tasks but struggles with complex number sequences and operations. The error distribution reveals key insights into the model’s strengths and weaknesses, guiding future improvements. A detailed error analysis for each category is included in Appendix  A.3.
@@ 5 Related works
Texual and Multimodal Reasoning dataset. There exist multiple datasets to test domain specific textual (math, science, medicine) QA and reasoning abilities of LLMs and VLMs knowledge such as SciBench Wang et al. (2023b), SciEvalSun et al. (2024), MMMU Yue et al. (2024), MathVista Lu et al. (2023), JEEbench Arora et al. (2023), MathVerse Zhang et al. (2024a), OlympiadBench He et al. (2024) and many others on based on real world images or other domains He et al. (2020); Soni et al. (2022); Liu et al. (2023); Thrush et al. (2022); Li et al. (2023). Even most existing visual and multi-modal reasoning datasets in the literature are domain-specific M3Exam Zhang et al. (2024b), RAVEN Zhang et al. (2019), or they involve reasoning about real-world images Liu et al. (2023); Wang et al. (2024); Thrush et al. (2022); Li et al. (2023) with basic commonsense reasoning questions such as CLEVR Johnson et al. (2017). However, current research has not thoroughly explored the capabilities of large models in addressing cognitive/critical reasoning problems for both textual and multi-modal data. NTSEBench is different from all already existing datasets in the literature because it explicitly focuses on testing cognitive reasoning abilities of large deep learning models. While benchmarks like BBH Suzgun et al. (2022) focus on text-based logical deduction and direction sense, NTSEBench offers a more holistic assessment by integrating both textual and visual elements. For example, it includes categories that combine text with visual answers, such as Venn diagrams and figure partition questions, which BBH does not cover. NTSEBench includes areas overlooked by other benchmarks, such as series, coding-decoding, and blood relations. With 2,728 multiple-choice questions, it provides a broader evaluation than the 770 puzzles in BBH.
@@
Furthermore, benchmarks like Jiang et al. (2024b) emphasize visual abstraction and reasoning but lack the integration of text and visual elements found in NTSEBench. Our dataset also goes beyond Raven’s Progressive Matrices (RPMs) Zhang et al. (2019) and related visual analogy problems, which are central to Małkiński and Mańdziuk (2023), by incorporating a wider array of cognitive tasks. This includes diverse categories such as coding-decoding, number and ranking, blood relations, and mathematical operations. NTSEBench provides a fine-grained analysis of model performance across 26 distinct problem categories and investigates various modeling strategies, setting it apart from other benchmarks. Datasets like Wang et al. (2024); Liu et al. (2023); Masry et al. (2022) explore mathematical reasoning in visual contexts but do not permit multiple images in the same prompt or reasoning across images. This is crucial for identifying patterns in categories of NTSEBench, such as Paper Folding and Cutting, Embedded Figure, Figure Partition, and Mirror/Water images.
@@
While ARC Chollet (2019) also targets general human cognitive understanding, its questions are primarily focused on visual tasks such as pattern completion, interpolation, denoising, and simple object counting. These tasks are relatively straightforward and often lack the need for complex reasoning. In contrast, NTSEBench spans 26 cognitive skills and integrates both textual and visual elements. By leveraging NTSE exam questions, it offers real-world relevance and facilitates quantitative evaluation through a multiple-choice format. Additionally, NTSEBench enables fine-grained performance analysis across cognitive dimensions, emphasizing knowledge integration and reasoning, which more closely mirrors human cognitive processes. These attributes make NTSEBench a more comprehensive and effective tool for evaluating AI’s cognitive capabilities.
@@
Language Writ Large Harnad (2024) delves into the theoretical foundations of large language models (LLMs) like ChatGPT, offering insights into their unexpected capabilities through the analysis of technology dialogues. In contrast, NTSEBench provides quantitative performance metrics for LLMs across a wide range of reasoning tasks. While LogicVista focuses on logical reasoning in visual contexts through five distinct tasks, NTSEBench covers a broader spectrum, spanning 26 categories that include both visual and textual reasoning. MMMU, on the other hand, spans six disciplines and 30 college subjects, providing a broad overview of subject-specific knowledge. In contrast, NTSEBench offers a deeper, more focused evaluation of cognitive reasoning skills, making it an essential tool for assessing the reasoning capabilities of LLMs.
@@
Zero shot and Few shot prompting engineering for textual and multi-modal input. Our work is also related to prompting design and prompt engineering for LLMs Brown et al. (2020); Chen et al. (2023); Gupta et al. (2023); Khot et al. (2022); Wei et al. (2022); Sahoo et al. (2024); Ali et al. (2024) and VLMs Xu et al. (2024); Bai et al. (2023); Liu et al. (2024); Dai et al. (2024).There are numerous studies on multi-modal and vision Chain-of-Thought (CoT) prompting Zhang et al. (2023); Shao et al. (2024). These studies overlook the ability of state-of-the-art vision-language models (VLMs) to process multiple images simultaneously, a limitation of older models with smaller context windows. Newer models can handle few-shot examples for CoT prompting, and we explore three prompting strategies enabled by their extended context windows.
@@ 6 Conclusion and Future Work
We developed a new dataset, NTSEBench, to assess the advanced analytical and logical reasoning capabilities of large deep learning models (LLMs and VLMs). We also propose four distinct modeling strategies for handling multi-modal data (text and images) across different question types in NTSEBench. These strategies enable us to conduct a fair and comprehensive comparison between closed and open-source models using both zero-shot and few-shot scenarios. Our findings show that both LLMs and VLMs struggle with advanced visual reasoning tasks, with VLMs performing worse on multi-modal questions than LLMs on textual ones. Proprietary models also consistently outperform open-source models, correctly predicting twice as many questions. Overall, our results underscore that NTSEBench poses significantly greater challenges for state-of-the-art LLMs and VLMs. Future directions. (a) Our results indicated that VLM models have difficulty predicting novel patterns, implying that addressing this challenge may involve either architectural modifications or the integration of generative models alongside VLMs for these question types.
(b) Given the limited data available for cognitive reasoning questions, we plan to use data augmentation strategies to increase the number of samples.
@@
While our dataset is new and sourced for different sources when compared to dataset already present in the literature, there still might be some overlap in reasoning questions, especially in textual reasoning. So all the dataset instances might not be independent and exclusive. This dataset is solely created in English, so no other languages are represented, therefore we cannot analyze whether language variations can have a significant impact on the reasoning capabilities of these large models. Our modelling strategies were limited to zero shot and few shot COT prompting. We did not evaluate whether fine-tuning these large models on a few examples from each of the categories would further improve results. This was due to limitation of both GPU resources and large cost of fine-tuning for Proprietary models. Finally, our human evaluation study involved only three annotators with undergraduate degrees, which may limit the generalizability of the results to the broader human population. We plan to address all these limitations in future extension of this work.
@@
As authors of this work, we affirm that our research adheres to the highest ethical standards in both its conduct and publication. We have carefully considered and addressed various ethical considerations inherent in computational linguistics methodologies to ensure responsible and fair practices. We prioritize transparency by providing detailed information to facilitate the reproducibility of our results. This includes sharing our code and datasets, which are sourced from publicly available datasets and handled in compliance with ethical guidelines established by the original authors. While our paper accurately reflects our experimental findings, we acknowledge the inherent variability associated with large language models, which we mitigate by maintaining a fixed temperature setting. We provide comprehensive descriptions of our annotations, dataset splits, models used, and prompting methods employed, ensuring the reproducibility of our work. For grammar correction, we utilize AI-based writing assistants, and for coding, we leverage tools like Copilot. Importantly, the genesis of our research ideas and the execution of our study were entirely independent of AI assistance.
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-23-1-2364. We extend our gratitude to the annotators who verified our flowcharts and corresponding question answer pairs.
@@ Appendix A Appendix
A.1 The Global Relevance of the NTSE Exam for AI
The NTSE exam, despite being conducted in India, holds significant relevance for the global AI community due to its unique focus on cognitive reasoning abilities rather than domain-specific knowledge. The NTSE’s diverse question categories as described in table  4, which assess various cognitive dimensions, offer a robust framework for evaluating AI models’ capacity to process information, recognize patterns, and solve problems across different domains. This emphasis on cognitive skills aligns with the pursuit of Artificial General Intelligence (AGI) , making the NTSE exam’s insights and challenges applicable to AI research and development on a global scale.
A.2 Additional Dataset Details
Table 1, reveals a skewed distribution across various question categories. Notably, textual categories like the Alphabet Test (ALP) and Mathematical Operations (MTO) contain 94 and 99 examples, respectively. In contrast, many vision-based categories are more challenging and typically include between 80 and 100 examples. For instance, the Non-Verbal Analogy (NVA) category, one of the most difficult, comprises 100 examples. While this skewed distribution could impact model performance, exploring its effects is beyond the scope of this manuscript and is left for future work.
@@ A.2 Additional Dataset Details
Modality Variations.
Since NTSEBench has multi-modal questions, options and solutions, we have results in eight different combination of modality type that can occur for question-options-solution triplet. Table 5 shows the count of each triplet option. NTSEBench has 1199 textual questions and remaining 1529 are multimodal questions.
Question Options Solutions # Samples
1199
381
70
18
330
126
403
201
A.2.1 Data Extraction Pipeline
The questions are extracted from the sources listed in section 2.1, with human intervention to monitor and rectify mistakes made by the automated pipeline. The data extraction pipeline involves first processing the PDF through MathPix OCR [^1] MathPix OCR to generate a Word file, which was then manually corrected for any errors. Next, we used the docxlatex 222docxlatex library to convert all equations into LaTeX expressions. Finally, we leveraged the PyMuPDF  [^3] PyMuPDF library to extract all text and images, extracting (1) Textual data i.e. Direction (Extra guidance on the context of the question), the Question, the correct Answer option and the Solution. (2) Vision Based data i.e. relevant image/s which we segregate into Direction images, Problem images, Option images, and Solution images. A total of 2,728 MCQ (multiple-choice questions) consisting of a total of 4,642 images across 26 categories of questions are extracted and NTSEBench is created along with the necessary metadata.
A.2.2 Dataset Sources Links
@@
A.3 Error Analysis for different question categories
@@
A.4 Additional Results
A.4.1 Preliminary Human Evaluation

### v3

@@ 4 Results and Analysis
NTSEBench presents a challenging task for SOTA LLMs and VLMs. Based on the findings in Tables 3 and 4, it is evident that the proposed dataset presents a challenging task for all state-of-the-art LLM and VLM models. None of the open-source models achieve accuracy exceeding 50% on text-only questions and 35% on multimodal questions, with proprietary models achieving 62% and 42% accuracy, respectively. Although o1-preview is categorized in a separately, it got more than 80% accuracy on text-only data, it cannot do multimodal reasoning. Many of the models tested did not even reach random selection/guess baseline of 24.52% (261 question had 5 options and 2467 question has 4 options), indicating that current LLMs and VLMs struggle with cognitive reasoning questions.
@@
Options Bias Ablation. From Gemini 1.5 Pro Results in Table  8 and  9 we can see that aggregate performance for both multimodal and text question is affected by the position of the correct option (variation in performance ranging from -4 to +6 percent when compared to random for text and -5 to +5 percent for multimodal categories). We can make a similar observation on GPT-4o (Tables  10 and  11); however, GPT-4o has smaller variation in performance than Gemini Pro. These variations suggest that the model’s responses to logical questions may be either memorized or guessed, indicating a potential limitation in LLMs and VLMs. Even o1-preview, as shown in Table 12, exhibits performance variations akin to those of GPT-4o, suggesting that advanced reasoning models are similarly affected by bias.
Error Analysis. We manually conducted an error analysis of 260 questions (10 from each question category) for Gemini 1.5 pro, and we identified distinct patterns in reasoning and error categorization. We have categorised errors based on the cognitive dimensions outlined in section 2. The Sankey diagram in Figure 2 illustrates how errors across various question categories correspond to specific error types. A key observation is that many errors arise from Pattern Recognition failures, especially in categories like Alphabet Tests, Non-verbal Analogy, and Series questions, where the model struggled with recurring patterns and sequence shifts, highlighting challenges in complex pattern-based reasoning. We also noted frequent errors in Spatial Reasoning and Logical Deduction tasks, particularly in spatial or diagrammatic questions such as Cube and Dice, Embedded Figure, and Paper Folding and Cutting. These questions often require pattern recognition, shape manipulation, or deducing logical relations from limited visual data. The figure also shows that errors in Quantitative Analysis were common in numerical questions such as Time and Clock and Mathematical Operations, indicating the model excels in simpler tasks but struggles with complex number sequences and operations. The error distribution reveals key insights into the model’s strengths and weaknesses, guiding future improvements. A detailed error analysis for each category is included in Appendix  A.2.
@@ 5 Related works
Texual and Multimodal Reasoning datasets. There exist multiple datasets to test domain specific textual (math, science, medicine) QA and reasoning abilities of LLMs and VLMs knowledge such as SciBench Wang et al. (2023b), SciEvalSun et al. (2024), MMMU Yue et al. (2024), MathVista Lu et al. (2023), JEEbench Arora et al. (2023), MathVerse Zhang et al. (2024a), OlympiadBench He et al. (2024) and many others based on real-world images or other domains He et al. (2020); Soni et al. (2022); Liu et al. (2023); Thrush et al. (2022); Li et al. (2023). Even most existing visual and multimodal reasoning datasets in the literature are domain-specific M3Exam Zhang et al. (2024b), RAVEN Zhang et al. (2019), or they involve reasoning about real-world images Liu et al. (2023); Wang et al. (2024); Thrush et al. (2022); Li et al. (2023) with basic common sense reasoning questions such as CLEVR Johnson et al. (2017). However, current research has not thoroughly explored the capabilities of large models in addressing cognitive/critical reasoning problems for both textual and multimodal data. NTSEBench is different from all already existing datasets in the literature because it explicitly focuses on testing cognitive reasoning abilities of large deep learning models. Although benchmarks like BBH Suzgun et al. (2022) focus on text-based logical deduction and direction sense, NTSEBench offers a more holistic assessment by integrating both textual and visual elements. For example, it includes categories that combine text with visual answers, such as Venn diagrams and figure partition questions, which BBH does not cover. NTSEBench also includes areas overlooked by other benchmarks, such as series, coding-decoding, and blood relations. With 2,728 multiple-choice questions, it provides a broader evaluation than the 770 puzzles in BBH.
@@
Furthermore, benchmarks such as Jiang et al. (2024b) emphasize visual abstraction and reasoning but lack the integration of text and visual elements found in NTSEBench. Our dataset also goes beyond Raven’s Progressive Matrices (RPMs) Zhang et al. (2019) and related visual analogy problems, which are central to Małkiński and Mańdziuk (2023), by incorporating a wider array of cognitive tasks. This includes diverse categories such as coding-decoding, number and ranking, blood relations, and mathematical operations. NTSEBench provides a fine-grained analysis of model performance across 26 distinct problem categories and investigates various modeling strategies, setting it apart from other benchmarks. Datasets such as Wang et al. (2024); Liu et al. (2023); Masry et al. (2022) explore mathematical reasoning in visual contexts but do not permit multiple images in the same prompt or reasoning across images. This is crucial for identifying patterns in categories of NTSEBench, such as Paper Folding and Cutting, Embedded Figure, Figure Partition, and Mirror/Water images.
@@
Although ARC Chollet (2019) also targets general human cognitive understanding, its questions are primarily focused on visual tasks such as pattern completion, interpolation, denoising, and simple object counting. These tasks are relatively straightforward and often lack the need for complex reasoning. By leveraging NTSE exam questions, our dataset offers real-world relevance and facilitates quantitative evaluation through a multiple-choice format. Additionally, NTSEBench enables fine-grained performance analysis across cognitive dimensions, emphasizing knowledge integration and reasoning, which more closely mirrors human cognitive processes. These attributes make NTSEBench a more comprehensive and effective tool for evaluating AI’s cognitive capabilities.
@@
Language Writ Large Harnad (2024) delves into the theoretical foundations of large language models (LLMs) like ChatGPT, offering insights into their unexpected capabilities through the analysis of technology dialogues. In contrast, NTSEBench provides quantitative performance metrics for LLMs across a wide range of reasoning tasks. MMMU Yue et al. (2024), on the other hand, spans six disciplines and 30 college subjects, providing a broad overview of subject-specific knowledge. In contrast, NTSEBench offers a deeper, more focused evaluation of cognitive reasoning skills, making it an essential tool for assessing the reasoning capabilities of LLMs.
@@
Zero shot and few shot prompt engineering for textual and multimodal input. Our work is also related to prompting design and prompt engineering for LLMs Brown et al. (2020); Chen et al. (2023); Gupta et al. (2023); Khot et al. (2022); Wei et al. (2022); Sahoo et al. (2024); Ali et al. (2024) and VLMs Xu et al. (2024); Bai et al. (2023); Liu et al. (2024); Dai et al. (2024). There are numerous studies on multimodal and vision Chain-of-Thought (CoT) prompting Zhang et al. (2023); Shao et al. (2024). These studies overlook the ability of state-of-the-art vision-language models (VLMs) to process multiple images simultaneously, a limitation of older models with smaller context windows. Newer models can handle few-shot examples for CoT prompting, and we explore three prompting strategies enabled by their extended context windows.
@@ 6 Conclusion and Future Work
We developed a new dataset, NTSEBench, to assess the advanced analytical and logical reasoning capabilities of large deep learning models (LLMs and VLMs). We also propose four distinct modeling strategies for handling multimodal data (text and images) across different question types in NTSEBench. These strategies enable us to conduct a fair and comprehensive comparison between proprietary and open-source models using both zero-shot and few-shot scenarios. Our findings show that both LLMs and VLMs struggle with advanced visual reasoning tasks, with VLMs performing worse on multi-modal questions than LLMs on textual ones. Proprietary models also consistently outperform open-source models, correctly predicting twice as many questions. Overall, our results underscore that NTSEBench poses significantly greater challenges for state-of-the-art LLMs and VLMs. Future directions. (a) Our results indicated that VLM models have difficulty predicting novel patterns, implying that addressing this challenge may involve either architectural modifications or the integration of generative models alongside VLMs for these question types. (b) Given the limited data available for cognitive reasoning questions, we plan to use data augmentation strategies to increase the number of samples.
@@
While our dataset is new and sourced for different sources when compared to datasets already present in the literature, there still might be some overlap in reasoning questions, especially in textual reasoning. So all the dataset instances might not be independent and exclusive. This dataset is solely created in English, so no other languages are represented; therefore, we cannot analyze whether language variations can have a significant impact on the reasoning capabilities of these large models. Our modelling strategies were limited to zero-shot and few-shot COT prompting. We did not evaluate whether fine-tuning these large models on a few examples from each of the categories would further improve results. This was due to the limitation of both GPU resources and large cost of fine-tuning for proprietary models. Finally, our human evaluation study involved only three annotators with undergraduate degrees, which may limit the generalizability of the results to the broader human population. We plan to address all these limitations in the future extension of this work.
@@
As authors of this work, we affirm that our research adheres to the highest ethical standards in both its conduct and publication. We have carefully considered and addressed various ethical considerations inherent in computational linguistics methodologies to ensure responsible and fair practices. We prioritize transparency by providing detailed information to facilitate the reproducibility of our results. This includes sharing our code and datasets, which are sourced from publicly available datasets and handled in compliance with ethical guidelines established by the original authors. Although our paper accurately reflects our experimental findings, we acknowledge the inherent variability associated with large language models, which we mitigate by maintaining a fixed temperature setting. We provide comprehensive descriptions of our annotations, dataset splits, models used, and prompting methods employed, ensuring the reproducibility of our work. For grammar correction, we utilize AI-based writing assistants, and for coding, we leverage tools such as Copilot. Importantly, the genesis of our research ideas and the execution of our study were entirely independent of AI assistance.
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-23-1-2364. We extend our gratitude to the annotators who verified our data extraction and corresponding question answer pairs. We extend our sincere appreciation to Jennifer Sheffield from the University of Pennsylvania for her administrative support. Lastly, we extend our appreciation to the reviewing team for their insightful comments.
@@ Appendix A Appendix
A.1 Additional Dataset Details
Table 1, reveals a skewed distribution across various question categories. Notably, textual categories like the Alphabet Test (ALP) and Mathematical Operations (MTO) contain 94 and 99 examples, respectively. In contrast, many vision-based categories are more challenging and typically include between 80 and 100 examples. For instance, the Non-Verbal Analogy (NVA) category, one of the most difficult, comprises 100 examples. Although this skewed distribution could impact model performance, exploring its effects is beyond the scope of this manuscript and is left for future work.
@@ A.1 Additional Dataset Details
A.1.1 Data Extraction Pipeline
The questions are extracted from the sources listed in section 2.1, with human intervention to monitor and rectify mistakes made by the automated pipeline. The data extraction pipeline involves first processing the PDF through MathPix OCR [^1] MathPix OCR to generate a Word file, which was then manually corrected for any errors. Next, we used the docxlatex 222docxlatex library to convert all equations into LaTeX expressions. Finally, we leveraged the PyMuPDF  [^3] PyMuPDF library to extract all text and images, extracting :
  • (1) Textual data i.e. direction (extra guidance on the context of the question), the question, the correct answer option and the solution. Any errors in text extraction was rectified by human annotator.
  • (2) Vision Based data i.e. relevant image/s which we segregate into direction images, problem images, option images, and solution images. Questions associated with low-quality images were excluded. A human annotator assisted in identifying these low-quality images.
A total of 2,728 MCQ (multiple-choice questions) consisting of a total of 4,642 images across 26 categories of questions are extracted and NTSEBench is created along with the necessary metadata.
The correct answers to the questions are also officially provided by the NTSE exam organizers (under NCERT). The exam is renowned for its high quality, as the questions are typically designed by subject-matter experts. Any disputes or challenges regarding incorrect answers or flawed questions are meticulously reviewed by NCERT before the release of the final answer key, ensuring the accuracy and reliability of the solutions, ensuring good quality data.
A.1.2 Dataset Sources Links
@@
A.2 Error Analysis for different question categories
@@
A.3 Additional Results
A.3.1 Preliminary Human Evaluation